OpenStudy (anonymous):

f(x) = 4x + 6, g(x) = 2x2 Find (fg)(x). 8x3 + 12x2 2x2 + 4x + 6 8x + 12 8x2 + 12x

OpenStudy (aum):

(f o g)(x) = f(g(x)) f(x) = 4x + 6 f(g(x)) = 4g(x) + 6 Substitute g(x) = 2x^2 on the right hand side of the above equation and simplify.